Determine the critical value of Scheffe’s test if the given value of the problem is K= 5 and N = 20 with 5% level of significance


1
Expert's answer
2021-02-24T13:24:09-0500

Given values are-

"k=5, N=20"

From The table of f-test

The critical value of f-test at 5% is "f_z=2.9899"


Critical value of scheffe's test "=(k-1)\times" critical value for f test

"=(5-1)\times f_z\\=(4)\times 2.9899\\=11.9596"


Hence critical value is "11.9596" .